OpenStudy (anonymous):

A population of bacteria grows at a rate proportional to size. on initially have 10,000 bacteria after 10 days and consisted of 20 000 bacteria. What is bacterial populations after 25 days

OpenStudy (anonymous):

So suppose x is number of bacteria dx/dt=-kx where k is some constant, take x to other side integrate dx/x=-kdt so ln(C/x)=kt where C is initial(10000) substitute the given case after 10 days ln2=10k k=ln2/10 Now substitute 25 and find the answer.

OpenStudy (anonymous):

My bad i think i put a negative sign by mistake dy/dx=kt ln(x/C)=kt and ln2=k10 , k=ln2/10

OpenStudy (dumbcow):

x = 10,000e^kt 20,000 = 10,000e^10k -> k = ln(2)/10 x = 10,000e^(25*ln(2)/10) = 56,568.54

OpenStudy (anonymous):

how value is ln2/10?

OpenStudy (anonymous):

ln(x/10000)=kt where t is obviously time, at t=10,x=20000(given) ln(20000/10000)=k*10 so ln2=10k,k=ln2/10

OpenStudy (anonymous):

for this case,we can use the same method? at the beginning of 1988, world population was estimated at 5.9 billion. It is said that in 2020, the population will reach 7.9 billion. a. How one can predict b. By using your assumptions dijawaban A, after how long the world's population will be doubled

OpenStudy (dumbcow):

yes if you assume exponential growth

OpenStudy (anonymous):

The same applies only if rate of population change proportional to population

OpenStudy (anonymous):

what means of rate of population change proportional to population? give me explain

OpenStudy (anonymous):

It means instantaneous change of population( If you draw a graph of population vs time then instantaneous change is given by slope of tangent or dy/dx) is some constant * population.

OpenStudy (anonymous):

oke,i understand.now...thanks you myfriend...

OpenStudy (anonymous):

No problem.